OpenStudy (anonymous):

lid is tightly sealed on a flask containing 2.25L Ar at 22C and .977atm. If the flask is heated to 102C, what is the pressure in the flask ? 2.86 1.27 4.53 .769 1.73 answers in atm

OpenStudy (anonymous):

Again, use pV=nRT. To perform this exercise, they want you to find a constant that will work on both problems- number of moles. You find it out, using n=pV/RT= 0.977*2.25*10^-3/(22+273)*8.31=8.96*10^-7 mol. Onto the second problem, we will have p=nRT/V= 8.96*10^-7*8.31*(102+273)/2.25*10^-3=1.24atm. Nearest answer will be 1.27atm